# Gene-DML: Dual-Pathway Multi-Level Discrimination for Gene Expression Prediction from Histopathology Images

### v1

@@ Abstract
Accurately predicting gene expression from histopathology images offers a scalable and non-invasive approach to molecular profiling, with significant implications for precision medicine and computational pathology. However, existing methods often underutilize the cross-modal representation alignment between histopathology images and gene expression profiles across multiple representational levels, thereby limiting their prediction performance. To address this, we propose Gene-DML, a unified framework that structures latent space through Dual-pathway Multi-Level discrimination to enhance correspondence between morphological and transcriptional modalities. The multi-scale instance-level discrimination pathway aligns hierarchical histopathology representations extracted at local, neighbor, and global levels with gene expression profiles, capturing scale-aware morphological-transcriptional relationships. In parallel, the cross-level instance-group discrimination pathway enforces structural consistency between individual (image/gene) instances and modality-crossed (gene/image, respectively) groups, strengthening the alignment across modalities. By jointly modeling fine-grained and structural-level discrimination, Gene-DML is able to learn robust cross-modal representations, enhancing both predictive accuracy and generalization across diverse biological contexts. Extensive experiments on public spatial transcriptomics datasets demonstrate that Gene-DML achieves state-of-the-art performance in gene expression prediction. The code and checkpoints will be released soon.
@@ 1 Introduction
Spatial transcriptomics (ST) has emerged as a transformative technology that enables high-resolution mapping of quantification of mRNA expression (gene expression) within intact tissue architectures [25, 24, 32]. Unlike bulk RNA sequencing, single-cell RNA sequencing (scRNA-seq) or single-nucleus RNA sequencing (snRNA-seq), which lacks spatial localization, spatial transcriptomics preserves the anatomical and microenvironmental context of gene expressions, making it invaluable for providing unprecedented insights into the spatial organization of molecular processes [33]. Intuitively, ST technology can be seen as converging on a gene expression matrix that captures the transcriptome at every spot (that is, a pixel, a cell, or a group of cells) [24], which is divided from large-scale whole-slide images (WSIs). Traditional ST methods, such as Visium [25], MERFISH [3], seqFISH+ [8], STARmap [30] et al. [9], tend to be time-consuming and expensive, relying on specialized instruments and extensive domain knowledge. This raises a critical question: how can we leverage recent advances in deep learning to accurately infer spatial gene expression directly from WSIs?
@@
Several recent studies have made substantial progress in gene expression prediction from WSIs [19, 35]. Unimodal approaches [10, 21, 37, 34, 6] typically employ supervised learning to predict gene profiles from WSI tiles, focusing on improving feature extraction to learn strong visual representations. However, despite the use of powerful image encoders, these unimodal methods often fail to incorporate gene-specific morphological information. The absence of alignment with gene expression weakens the model’s ability to capture meaningful biological signals, often resulting in overfitting to high-dimensional gene expression counts and weak generalization across patient samples.
@@
To explore image-gene relationships more effectively, bimodal methods [33, 20, 13, 18, 9] co-embed WSI tiles and gene expression profiles in a shared space. While these approaches enable cross-modal learning, they typically model only single-level, pairwise relationships between individual tiles and gene instances. By treating each image-gene pair independently, they fail to capture the hierarchical alignment information inherent in neither WSI tile nor gene expression profile instances, triggering overfitting to the instance-level noise. Thus, multi-level alignments between image and gene modalities are essential for constructing a robust and expressive latent space that faithfully encodes both WSI morphology and gene expression; otherwise, neglecting this limits model generalization and impairs the learning of biologically coherent representations.
@@
To this end, we propose Gene-DML, a novel framework that enhances Gene-image correspondence by performing Dual-pathway Multi-Level discrimination. It jointly captures the scale-aware morphological-transcriptional relationship between multi-scaled WSI tiles and gene expression profiles at the instance level and instance-group level, that is, WSI tile instances aligned with clustered gene-expression groups and vice versa. By jointly modeling fine-grained and high-level discrimination, Gene-DML enables the learning of robust cross-modal representations, enhancing both predictive accuracy and generalization across diverse biological contexts.
Our contributions are summarized as follows:
@@ 2.1 Gene Expression Prediction
Early methods for spatial gene expression prediction from histology, such as ST-Net [10], employed a transfer learning strategy by fine-tuning a DenseNet121 model [12] pre-trained on ImageNet to capture image features and map WSI tiles to gene expression labels. Building upon this, HisToGene [21] integrates Vision Transformers (ViT) to capture global contextual dependencies between image tiles within a WSI. Further advancing this direction, Hist2ST [37] incorporates ConvMixer [27] and a graph convolution network (GCN) post-initially for enhanced image tile embeddings by aggregating neighborhood information. Recent multi-scale models, TRIPLEX [6] and M2OST [29] improved performance by integrating spot-, neighborhood-, and slide-level features. These architectures leveraged hierarchical modeling and fusion techniques to capture both fine-grained morphology and global tissue context. EGN [34] is an alternative paradigm in which image embedding is optimized by selecting the most analogous exemplars from the target spot within a WSI to enhance gene expression prediction. Despite these advances, these unimodal methods lack semantic alignment with gene expression and struggle to generalize across samples. Bimodal contrastive learning has gained traction. To explore the image-gene relationship for more scalable and broader potential approaches, bimodal models, BLEEP [33] and MclSTExp [20] with Transformer encoders align image and gene embeddings through contrastive loss, enabling gene querying via latent similarity. RankByGene [13] and ST-Align [18] further improved alignment through ranking consistency and large-scale pretraining, while Stem [38] explored generative modeling using diffusion techniques. Existing contrastive or generative methods focus mainly on pairwise alignment and overlook deeper morphological-transcriptional correspondences across image scales, which is essential for biological semantic understanding.
@@ 2.2 Multi-scaled Image Feature Extraction
Several recent works have explored multi-scale modeling for WSIs to capture both fine-grained and contextual features. HIPT by Chen et al. [4] employs hierarchical self-supervised learning to learn patch-level and region-level features via contrastive objectives, enabling effective representation learning across image resolutions. TRIPLEX [6] is proposed by integrating multi-resolution features from target, neighbor, and global tissue levels using a supervised fusion strategy. M2OST [29] leverages multi-scale WSI patches with Transformer-based many-to-one regression for accurate gene prediction. Cross-scale MIL architectures have also shown promise. MEGT [7] and DSF-WSI [28] adopt dual-branch architectures with separate high- and low-resolution encoders, fusing outputs using graph‑Transformer modules or masked‐jigsaw tasks. Despite these advances, existing methods suffer from the following limitations: (1) most focus exclusively on unimodal representation learning without leveraging alignment of transcriptomic modality data, and (2) fusion-based designs often lack contrastive objectives that explicitly enforce scale-consistent representation. In contrast, Gene-DML incorporates multi-scale instance-level discrimination that explicitly aligns local, neighbor, and global image features with gene expression, enriching cross-modal representations and enhancing semantic consistency across scales.
@@ 2.3 Cross-Level Instance-Group Discrimination
Cross-Level Discrimination (CLD) [31] is proposed to improve instance-level contrastive learning by incorporating dynamic group-level structure within a single modality. By pairing each instance with its group centroid, CLD mitigates over-separation of semantically similar samples and improves representation stability. Subsequent work has adapted instance-group techniques within a single modality and for video-text alignment, e.g., CrossCLR [39], TCL [17], by incorporating grouping or de-biasing false negatives, these methods empirically show improved retrieval, classification, and stability [17]. However, they remain unimodal or restricted to intra-modality alignment. In medical imaging and genomics, multimodal frameworks like ContIG [26] and MGI [17] have adopted instance-level contrastive learning across modalities (e.g., image-genetic or image-omics pairs), but these lack mechanisms to model group-level relationships or latent clustering in cross-modal settings. They typically use pairwise contrastive objectives akin to CLIP [23], which may overemphasize instance-level uniqueness and discount shared biological patterns. To bridge these gaps, Gene-DML integrates a cross-level instance-group discrimination mechanism within two modalities: images and transcriptomics. It aligns each image instance with gene-expression groups and vice versa, ensuring the model learns shared semantic structure between morphology and transcriptional activity. This bidirectional group alignment improves robustness to noise, enhances positive/negative contrastive signal, and supports generalization to unseen tissue domains.
@@ 3.1.1 Multi-scale Image Instance Feature
A single spot’s gene expression is influenced not only by its WSI tile structure but also by its surrounding environment [6]. Each WSI is split into tiles centered on spatial transcriptomics spots, extracted at three spatial scales: local tiles $I_{L}$ focus on local-level detail with a size of $224\times 224$, neighbor tiles $I_{N}$ capture microenvironment context with a size of $1120\times 1120$, and the global scale $I_{G}$ represents the WSI’s global tissue structure. These tiles are encoded via a shared Multi-Scale Image Encoder. Specifically, we utilize the UNI [5] model, pretrained using more than 100 million H&E-stained WSIs, for histopathology image feature extraction. In order to avoid the overfitting problem, the base feature extraction model UNI is kept frozen to ensure the generalization of Gene-DML. The features from all three scales are individually projected into $512$ dimensions via separate linear layers.
@@
To efficiently capture global spatial context in WSIs without incurring high computational costs, inspired by [6], the local-scale spot features are first extracted by the UNI model, then passed into a transformer block.
@@ 3.1.2 Gene Expression Instance Feature
where $i=1,...,n$. Here, GeneEncoder first maps raw gene profiles into an intermediate embedding space of dimension $d$, capturing high-level transcriptomic structure, consisting of two linear layers with GELU [11] activation and dropout for regularization. This is followed by a FeedForward (FFN) network consisting of two linear layers with a non-linear activation and dropout, which refines the representation and projects it into the shared latent space. This design allows Gene-DML to learn instance-level gene features that align with corresponding histology-derived representations.
@@ 3.2.1 Feature Grouping
To structurally organize the latent space and uncover shared biological semantics, we perform modality-specific feature grouping based on the group features $\{I^{E_{\text{clu}}}_{i}\}$ and $\{G^{E_{\text{clu}}}_{i}\}$. Specifically, the $K$-means strategy [36] is applied to unit-normalized vectors to form $k$ local clusters in each modality. This yields sets of centroids $\{C^{I}_{1},\dots,C^{I}_{k}\}$ and $\{C^{G}_{1},\dots,C^{G}_{k}\}$, which are $\{C_{i}\}$ in gradient-colored circles and $\{C_{i}\}$ in red circles, respectively, as shown in Fig. 1.
@@ 3.3 Loss Component
Our method follows the CLIP [22] framework that maximizes the cosine similarity of the image and gene embeddings of the $N$ real pairs while minimizing the cosine similarity of the embeddings of the $N^{2}-N$ incorrect pairings at the image-gene instance level. At the multi-scale instance-level, the contrastive loss is applied on WSI embeddings from different scales with gene embedding. At the instance-group level, Gene-DML applies a batch-wise contrastive loss across views between group features $I^{E_{\text{ins}}}_{i}$ and group centroids $C^{G}$, and vice versa for $G^{E_{\text{ins}}}_{i}$ (with $C^{I}$), to supervise the consistency of grouping relationships.
@@ 3.3.1 Multi-scale Instance-level Loss
Inspired by the contrastive loss used in [33], Gene-DML computes the internal multi-scaled similarities between the image or the gene expression features in advance, denoted as $\mathbf{T}$, increasing the coherence of the latent space for joint embedding. Then, cross-entropy loss is applied to align the multi-scaled image and gene expression instance features:
@@ 4 Experiments and Results
ST-Net [10] 20’ $0.260\pm 0.04$ $0.194\pm 0.11$ $0.345\pm 0.16$ $0.209\pm 0.02$ $0.116\pm 0.06$ $0.223\pm 0.10$ $0.294\pm 0.07$ $0.274\pm 0.08$ $0.382\pm 0.08$
HistoGene [21] 21’ $0.314\pm 0.09$ $0.168\pm 0.12$ $0.302\pm 0.19$ $0.194\pm 0.05$ $0.100\pm 0.05$ $0.219\pm 0.12$ $0.270\pm 0.09$ $0.133\pm 0.06$ $0.261\pm 0.13$
Hist2ST [37] 22’ $0.285\pm 0.08$ $0.118\pm 0.10$ $0.248\pm 0.17$ $\underline{0.181}\pm 0.02$ $0.044\pm 0.02$ $0.099\pm 0.03$ $1.291\pm 0.65$ $0.004\pm 0.01$ $0.053\pm 0.01$
EGN [34] 23’ $0.241\pm 0.06$ $0.197\pm 0.11$ $0.328\pm 0.17$ $0.192\pm 0.02$ $0.111\pm 0.05$ $0.203\pm 0.09$ $0.281\pm 0.08$ $0.281\pm 0.06$ $0.388\pm 0.06$
BLEEP [33] 23’ $0.277\pm 0.05$ $0.151\pm 0.11$ $0.277\pm 0.16$ $0.235\pm 0.02$ $0.095\pm 0.05$ $0.193\pm 0.10$ $0.297\pm 0.05$ $0.269\pm 0.07$ $0.396\pm 0.08$
TRIPLEX [6] 24’ $\underline{0.228}\pm 0.07$ $\underline{0.314}\pm 0.14$ $\underline{0.497}\pm 0.17$ $0.202\pm 0.02$ $\underline{0.206}\pm 0.07$ $0.352\pm 0.10$ $\underline{0.268}\pm 0.09$ $\underline{0.374}\pm 0.07$ $\underline{0.490}\pm 0.07$
M2OST [29] 25’ $0.302\pm 0.06$ $0.231\pm 0.10$ $0.410\pm 0.17$ $0.278\pm 0.03$ $0.201\pm 0.07$ $\underline{0.353}\pm 0.10$ $0.271\pm 0.02$ $0.300\pm 0.05$ $\underline{0.490}\pm 0.05$

### v3

@@ Abstract
Accurately predicting gene expression from histopathology images offers a scalable and non-invasive approach to molecular profiling, with significant implications for precision medicine and computational pathology. However, existing methods often underutilize the cross-modal representation alignment between histopathology images and gene expression profiles across multiple representational levels, thereby limiting their prediction performance. To address this, we propose Gene-DML, a unified framework that structures latent space through Dual-pathway Multi-Level discrimination to enhance correspondence between morphological and transcriptional modalities. The multi-scale instance-level discrimination pathway aligns hierarchical histopathology representations extracted at local, neighbor, and global levels with gene expression profiles, capturing scale-aware morphological-transcriptional relationships. In parallel, the cross-level instance-group discrimination pathway enforces structural consistency between individual (image/gene) instances and modality-crossed (gene/image, respectively) groups, strengthening the alignment across modalities. By jointly modeling fine-grained and structural-level discrimination, Gene-DML is able to learn robust cross-modal representations, enhancing both predictive accuracy and generalization across diverse biological contexts. Extensive experiments on public spatial transcriptomics datasets demonstrate that Gene-DML achieves state-of-the-art performance in gene expression prediction. The code and processed datasets are available at https://github.com/YXSong000/Gene-DML.
@@ 1 Introduction
Spatial transcriptomics (ST) has emerged as a transformative technology that enables high-resolution mapping of quantification of mRNA expression (gene expression) within intact tissue architectures [stahl_2021, anjali_2021, Wang_2024]. Unlike bulk RNA sequencing [wang2025mirror], single-cell RNA sequencing (scRNA-seq) or single-nucleus RNA sequencing (snRNA-seq), which lacks spatial localization, spatial transcriptomics preserves the anatomical and microenvironmental context of gene expressions, making it invaluable for providing unprecedented insights into the spatial organization of molecular processes [xie_2024]. Intuitively, ST technology can be seen as converging on a gene expression matrix that captures the transcriptome at every spot (that is, a pixel, a cell, or a group of cells) [anjali_2021], which is divided from large-scale whole-slide images (WSIs). Traditional ST methods, such as Visium [stahl_2021], MERFISH [chen2015spatially], seqFISH+ [Eng_2019], STARmap [wang2018three] et al[ganguly2025merge], tend to be time-consuming and expensive, relying on specialized instruments and extensive domain knowledge. This raises a critical question: how can we leverage recent advances in deep learning to accurately infer spatial gene expression directly from WSIs?
@@
Several recent studies have made substantial progress in gene expression prediction from WSIs [Mejia_2024, yang_2024]. Unimodal approaches [he_2020, pang_2021, zeng_2022, Yang_2023, chung_2024] typically employ supervised learning to predict gene profiles from WSI tiles, focusing on improving feature extraction to learn strong visual representations. However, despite the use of powerful image encoders, these unimodal methods often fail to incorporate gene-specific morphological information. The absence of alignment with gene expression weakens the model’s ability to capture meaningful biological signals, often resulting in overfitting to high-dimensional gene expression counts and weak generalization across patient samples.
@@
To explore image-gene relationships more effectively, bimodal methods [min_2024, huang_2024, lin_2024] co-embed WSI tiles and gene expression profiles in a shared space. While these approaches enable cross-modal learning, they typically model only single-level, pairwise relationships between individual tiles and gene instances. By treating each image-gene pair independently, they fail to capture the hierarchical alignment inherent in neither WSI tile nor gene expression profile instances, triggering overfitting to the instance-level noise. Thus, multi-level alignments between image and gene modalities are essential for constructing a robust and expressive latent space that faithfully encodes both histopathology morphology [10.1007/978-3-031-73464-9_17, 8363568, 9731234] and gene expression; otherwise, the model’s generalization and the learning of biologically coherent representations are compromised.
@@
To this end, we propose Gene-DML, a novel framework that enhances Gene-image correspondence by performing Dual-pathway Multi-Level discrimination. It jointly captures the scale-aware morphological-transcriptional relationship between multi-scaled WSI tiles and gene expression profiles at the instance level and instance-group level, that is, WSI tile instances aligned with clustered gene-expression groups and vice versa. By jointly modeling fine-grained and high-level discrimination, Gene-DML enables the learning of robust cross-modal representations, enhancing both predictive accuracy and generalization across diverse biological contexts. Contributions are summarized as:
@@ 2.1 Gene Expression Prediction
Early methods for spatial gene expression prediction from histology [fan2024learning], such as ST-Net [he_2020], employed a transfer learning strategy by fine-tuning a DenseNet121 model [huang_2017] pre-trained on ImageNet to capture image features and map WSI tiles to gene expression labels. Building upon this, HisToGene [pang_2021] integrates Vision Transformers (ViT) to capture global contextual dependencies between image tiles within a WSI. Further advancing this direction, Hist2ST [zeng_2022] incorporates ConvMixer [Trockman_2022] and a graph convolution network (GCN) post-initially for enhanced image tile embeddings by aggregating neighborhood information. Recent multi-scale models, TRIPLEX [chung_2024] and M2OST [wang2025m2ost] improved performance by integrating spot-, neighborhood-, and slide-level features. These architectures leveraged hierarchical modeling and fusion techniques to capture both fine-grained morphology and global tissue context. EGN [Yang_2023] is an alternative paradigm in which image embedding is optimized by selecting the most analogous exemplars from the target spot within a WSI to enhance gene expression prediction. Despite these advances, these unimodal methods lack semantic alignment with gene expression and struggle to generalize across samples. Bimodal contrastive learning has gained traction. To explore the image-gene relationship for more scalable and broader potential approaches, bimodal models, BLEEP [xie_2024] and MclSTExp [min_2024] with Transformer encoders align image and gene embeddings through contrastive loss, enabling gene querying via latent similarity. RankByGene [huang_2024] and ST-Align [lin_2024] further improved alignment through ranking consistency and large-scale pretraining, while Stem [zhu2025diffusion] explored generative modeling using diffusion techniques [wan2025tedvit]. Existing contrastive or generative methods focus mainly on pairwise alignment and overlook deeper morphological-transcriptional correspondences across image scales, which is essential for biological semantic understanding.
@@ 2.2 Multi-scaled Image Feature Extraction
Several recent works have explored multi-scale modeling for WSIs to capture both fine-grained and contextual features [fang2025scsamdebiasingmorphologydistributional]. HIPT by Chen et al[chen2022scaling] employs hierarchical self-supervised learning to learn patch-level and region-level features via contrastive objectives, enabling effective representation learning across image resolutions. TRIPLEX [chung_2024] is proposed by integrating multi-resolution features from target, neighbor, and global tissue levels using a supervised fusion strategy. M2OST [wang2025m2ost] leverages multi-scale WSI patches with Transformer-based many-to-one regression for accurate gene prediction. Cross-scale MIL architectures have also shown promise. MEGT [ding2023multi] and DSF-WSI [wang2023dual] adopt dual-branch architectures with separate high- and low-resolution encoders, fusing outputs using graph‑Transformer modules or masked‐jigsaw tasks. Despite these advances, existing methods suffer from the following limitations: (1) most focus exclusively on unimodal representation learning without leveraging alignment of transcriptomic modality data, and (2) fusion-based designs often lack contrastive objectives that explicitly enforce scale-consistent representation. In contrast, Gene-DML incorporates multi-scale instance-level discrimination that explicitly aligns local, neighbor, and global image features with gene expression, enriching cross-modal representations and enhancing semantic consistency across scales.
@@ 2.3 Cross-Level Instance-Group Discrimination
Cross-Level Discrimination (CLD) [wang_2021] is proposed to improve instance-level contrastive learning by incorporating dynamic group-level structure within a single modality. By pairing each instance with its group centroid, CLD mitigates over-separation of semantically similar samples and improves representation stability. Subsequent work has adapted instance-group techniques within a single modality and for video-text alignment, e.g., CrossCLR [zolfaghari2021crossclr], TCL [li2022twin], by incorporating grouping or de-biasing false negatives, these methods empirically show improved retrieval, classification, and stability [li2022twin]. However, they remain unimodal or restricted to intra-modality alignment. In medical imaging and genomics, multimodal frameworks like ContIG [taleb2022contig] and MGI [li2022twin] have adopted instance-level contrastive learning across modalities (e.g., image-genetic or image-omics pairs), but these lack mechanisms to model group-level relationships or latent clustering in cross-modal [chen2024gaussian]. They typically use pairwise contrastive objectives akin to CLIP [radford2021learning, wang2023mosaic], which may overemphasize instance-level uniqueness and discount shared biological patterns. To bridge these gaps, Gene-DML integrates a cross-level instance-group discrimination mechanism within two modalities: images and transcriptomics. It aligns each image instance with gene-expression groups and vice versa, ensuring the model learns shared semantic structure between morphology and transcriptional activity. This bidirectional group alignment improves robustness to noise, enhances positive/negative contrastive signal, and supports generalization to unseen tissue domains.
@@ 3.1.1 Multi-scale Image Instance Feature
A single spot’s gene expression is influenced not only by its WSI tile structure but also by its surrounding environment [chung_2024]. Each WSI is split into tiles centered on spatial transcriptomics spots, extracted at three spatial scales: local tiles $I_{L}$ focus on local-level detail, neighbor tiles $I_{N}$ capture microenvironment context, and the global scale $I_{G}$ represents the WSI’s global tissue structure. These tiles are encoded via a shared Multi-Scale Image Encoder. Specifically, we utilize the UNI [chen2024uni] model, pretrained using more than 100 million H&E-stained WSIs, for histopathology image feature extraction. In order to avoid the overfitting problem, the base feature extraction model UNI is kept frozen to ensure the generalization of Gene-DML. The features from all three scales are individually projected into $512$ dimensions via separate linear layers.
@@
To efficiently capture global spatial context in WSIs without incurring high computational costs, inspired by [chung_2024], the global-scale representation is concatenated with all local-scale tile features extracted by the pre-trained UNI model, which is then encoded through the transformer block for the ultimate global-scale token.
@@ 3.1.2 Gene Expression Instance Feature
where i=1,,ni=1,...,n. Here, GeneEncoder first maps raw gene profiles into an intermediate embedding space of dimension $d$, capturing high-level transcriptomic structure, consisting of two linear layers with GELU [hendrycks2016gaussian] activation and dropout for regularization. This is followed by a FeedForward (FFN) network consisting of two linear layers with a non-linear activation and dropout, which refines the representation and projects it into the shared latent space. This design allows Gene-DML to learn instance-level gene features that align with corresponding histology-derived representations.
@@ 3.2.1 Feature Grouping
To structurally organize the latent space and uncover shared biological semantics, we perform modality-specific feature grouping based on the group features $\{I^{E_{\text{clu}}}_{i}\}$ and $\{G^{E_{\text{clu}}}_{i}\}$. Specifically, the $K$-means strategy [Ye_2007] is applied to unit-normalized vectors to form $k$ local clusters in each modality. This yields sets of centroids $\{C^{I}_{1},\dots,C^{I}_{k}\}$ and $\{C^{G}_{1},\dots,C^{G}_{k}\}$, which are $\{C_{i}\}$ in gradient-colored circles and $\{C_{i}\}$ in red circles, respectively, as shown in Fig. 1.
@@ 3.3 Loss Component
Our method follows the CLIP [radford2021learning] framework that maximizes the cosine similarity of the image and gene embeddings of the $N$ real pairs while minimizing the cosine similarity of the embeddings of the $N^{2}-N$ incorrect pairings at the image-gene instance level. At the multi-scale instance-level, the contrastive loss is applied on WSI embeddings from different scales with gene embedding. At the instance-group level, Gene-DML applies a batch-wise contrastive loss across views between group features $I^{E_{\text{ins}}}_{i}$ and group centroids $C^{G}$, and vice versa for $G^{E_{\text{ins}}}_{i}$ (with $C^{I}$), to supervise the consistency of grouping relationships.
@@ 3.3.1 Multi-scale Instance-level Loss
Inspired by the contrastive loss used in [xie_2024], Gene-DML computes the internal multi-scaled similarities between the image or the gene expression features in advance, denoted as $\mathbf{T}$, increasing the coherence of the latent space for joint embedding. Then, cross-entropy loss is applied to align the multi-scaled image and gene expression instance features:
@@ 4 Experiments and Results
ST-Net [he_2020] 20’ $0.260\pm 0.04$ $0.194\pm 0.11$ $0.345\pm 0.16$ $0.209\pm 0.02$ $0.116\pm 0.06$ $0.223\pm 0.10$ $0.294\pm 0.07$ $0.274\pm 0.08$ $0.382\pm 0.08$
HistoGene [pang_2021] 21’ $0.314\pm 0.09$ $0.168\pm 0.12$ $0.302\pm 0.19$ $0.194\pm 0.05$ $0.100\pm 0.05$ $0.219\pm 0.12$ $0.270\pm 0.09$ $0.133\pm 0.06$ $0.261\pm 0.13$
Hist2ST [zeng_2022] 22’ $0.285\pm 0.08$ $0.118\pm 0.10$ $0.248\pm 0.17$ $\underline{0.181}\pm 0.02$ $044\pm 0.02 0.0$$099\pm 0.03$ $1.291\pm 0.65$ $004\pm 0.01 0.0$$053\pm 0.01$
EGN [Yang_2023] 23’ $0.241\pm 0.06$ $0.197\pm 0.11$ $0.328\pm 0.17$ $0.192\pm 0.02$ $0.111\pm 0.05$ $0.203\pm 0.09$ $0.281\pm 0.08$ $0.281\pm 0.06$ $0.388\pm 0.06$
BLEEP [xie_2024] 23’ $0.277\pm 0.05$ $0.151\pm 0.11$ $0.277\pm 0.16$ $235\pm 0.02 0.0$$095\pm 0.05$ $0.193\pm 0.10$ $0.297\pm 0.05$ $0.269\pm 0.07$ $0.396\pm 0.08$
TRIPLEX [chung_2024] 24’ $\underline{0.228}\pm 0.07$ $\underline{0.314}\pm 0.14$ $\underline{0.497}\pm 0.17$ $0.202\pm 0.02$ $\underline{0.206}\pm 0.07$ $0.352\pm 0.10$ $\underline{0.268}\pm 0.09$ $\underline{0.374}\pm 0.07$ $\underline{0.490}\pm 0.07$
M2OST [wang2025m2ost] 25’ $0.302\pm 0.06$ $0.231\pm 0.10$ $0.410\pm 0.17$ $0.278\pm 0.03$ $0.201\pm 0.07$ $\underline{0.353}\pm 0.10$ $0.271\pm 0.02$ $0.300\pm 0.05$ $\underline{0.490}\pm 0.05$